# Quantum Langevin Dynamics for Optimization

### v2

@@ Abstract
We initiate the study of utilizing Quantum Langevin Dynamics (QLD) to solve optimization problems, particularly those nonconvex objective functions that present substantial obstacles for traditional gradient descent algorithms. Specifically, we examine the dynamics of a system coupled with an infinite heat bath. This interaction induces both random quantum noise and a deterministic damping effect to the system, which nudge the system towards a steady state that hovers near the global minimum of objective functions. We theoretically prove the convergence of QLD in convex landscapes, demonstrating that the average energy of the system can approach zero in the low temperature limit with an exponential decay rate correlated with the evolution time. Numerically, we first show the energy dissipation capability of QLD by retracing its origins to spontaneous emission. Furthermore, we conduct detailed discussion of the impact of each parameter. Finally, based on the observations when comparing QLD with the classical Fokker-Plank-Smoluchowski equation, we propose a time-dependent QLD by setting temperature and $\hbar$ as time-dependent parameters, which can be theoretically proven to converge better than the time-independent case and also outperforms a series of state-of-the-art quantum and classical optimization algorithms in many nonconvex landscapes.
@@ 1 Introduction
$\dot{Y}=\frac{\mathrm{i}}{\hbar}[H_{{total}},Y].$ (2)
However, we are manipulating operators in quantum mechanics, which makes the analysis of above equation of motion highly challenging. Therefore, we adopt the Born-Markov approximation and use path integral approach [48, 8] to solve this challenge. When the correlation time of the noise is significantly shorter than the typical time scale of the damped motion of the system, we obtain the quantum Brownian motion master equation [32, 48, 8]:
@@
where $H=p^{2}/(2m)+V$, $p$ is the momentum operator of the system, $x$ is the position operator, and $\eta=\gamma/2m$ represents the characteristic damping rate with system mass $m$, $k$ is the Boltzmann constant. However, it is well known that this master equation is only applicable in high-temperature scenarios and fails to satisfy the positivity constraint of the density operator [29, 19, 44]. Ultimately, we refer to Ref. [29] and employ Lindblad functional (Eq. (14)) to describe the quantum Langevin dynamics, with a detailed introduction provided in Section 2.2.
@@
where $V(x)$ is assumed to be continuously differentiable in the domain and is also the potential energy of the system in Eq. (14). One can implement Quantum Langevin Dynamics (QLD) by starting from an easily prepared initial state and evolving the quantum system according to Eq. (14). The global minimum of the objective function $V(x)$ can be obtained by measuring the position observable $x$ at the end of the evolution. Therefore, our objective is to establish a theoretical foundation guaranteeing the convergence to the global minimum of QLD. When the parameters of QLD is time-independent, namely, $\eta,\hbar$ and $T$ are constants during evolution, we establish the convergence of QLD within convex landscapes and specific instances of slightly nonconvex scenarios. Informally speaking, with approximation of $\hbar\Omega\ll kT$ where $\Omega$ is the oscillator frequency of the ground state near the global minimum, we have the convergence rate of the average potential,
$\left\langle{V}\right\rangle_{t}-V(x^{*})\leq 2kf(T)+O(e^{-\eta t}),$ (5)
@@
where $\left\langle{V}\right\rangle_{t}$ denotes the expectation value of $V$ with respect to the density matrix of the system, $V(x^{*})$ is the potential value at the global minimum $x^{*}$ and $f(T)$ is a function with respect to $T$ that satisfies $\lim_{T\to 0}f(T)=0$. Consequently, the average energy of the system can approach zero in the low temperature limit. When $V$ is a quadratic form, the evolution of the system has an analytical solution. One interesting discovery related to the quadratic potential is that the system can reach the ground state of the quadratic potential when the evolution time $t\to+\infty$ and $T\to 0$. Formal description can be found in Section 3. However, it is always challenging to give a convergence in the general nonconvex case. Despite our best efforts, we are constrained to confine our theoretical discussions to slightly nonconvex cases. Thus, for other nonconvex landscapes, we resort to meticulous numerical experiments to compensate for the absence of theoretical assurances.
@@
In the realm of numerical experiments, we initially substantiate our intuition regarding the fundamental cause of energy dissipation in QLD. Noteworthy is the observation that despite the finite nature of the heat bath, the system retains the capacity to escape local minima, transitioning from excited states to ground states due to spontaneous emission. In alignment with our theoretical findings, we execute experiments within quadratic frameworks, drawing comparisons with analytic solutions to demonstrate a high degree of concordance. To offer optimal strategies for harnessing thermal and tunneling effects when confronting intricate nonconvex problems, we engage in a detailed discussion delineating the role of each parameter. Broadly speaking, our results indicate that $\eta$ determines the convergence rate, temperature $T$ reflects the thermal effect to some extent and $\hbar$ reflects the tunneling effect.
@@
Building upon these observations, we find it imperative to explore the realm of time-dependent QLD. Specifically, we recognize the need for pronounced thermal and tunneling effects initially to facilitate escape from local minima. As the system converges to its final stage, these effects need to be gradually deactivated. Consequently, we present a theoretical analysis for time-dependent QLD and propose a cooling strategy for implementation. The convergence of time-dependent QLD is similar to the time-independent one when $\hbar=\hbar(t)$ decreasing monotonically with respect to time and $\eta=\eta(t)$, which says $\left\langle{V}\right\rangle_{t}-V(x^{*})\leq 2kT+O(e^{-\int_{0}^{t}\eta(t^{\prime})\mathrm{d}t^{\prime}})$. For the case of time-dependent temperature, there exists a cooling strategy $T=T(t)$ that guarantees convergence. In numerical experiments, time-dependent QLD significantly outperforms its time-independent counterpart and surpasses other algorithms, including QHD, QAA, and SGD, across a diverse array of landscapes. In theory, the query complexity of QLD is $\tilde{O}(t)$, which is also among the top tier of all other algorithms.
The rest of this paper is laid out as follows. In Section 2, we introduce our notations and provide details for quantum Langevin equation and explicit description of QLD in Eq. (14). Theoretically, we prove the convergence of QLD in convex case (Theorem 1) and slightly nonconvex case (Theorem 3), refer to Section 3. Numerically, we first trace back to spontaneous emission to show the essential cause of energy dissipation in QLD, which verifies our intuition, see Section 4.1. Then we verify our theoretical results of quadratic case in Section 4.3 and discuss the roles of each parameter and their relations with thermal and tunneling effects in Section 4.4. In Section 5 we compare QLD with classical dynamics, which informs the development of time-dependent QLD to fully leverage the advantages inherent in quantum dynamics. In Section 5.1, we prove the convergence of time-dependent QLD in convex case. The performance of time-dependent QLD is evaluated in eight representative landscapes, demonstrating superior performance through comparisons with QHD, QAA, and classical algorithms, shown in Section 5.6. Finally, we evaluate and compare the query complexity of QLD with other algorithms in Appendix B, demonstrating that QLD is superior to other algorithms in terms of query complexity.
@@ 2 Preliminaries
Notations. We denote the Hamiltonian of system as $H$, the Hamiltonian of heat bath as $H_{B}$, the Hamiltonian of heat bath and system as $H_{\text{total}}=H+H_{B}$. Let $E_{k}$ be the kinetic energy of the system and $V$ be the potential energy of the system, so $H=E_{k}+V$. Given two quantum operators $A$ and $B$, we use $[A,B]=AB-BA$ for denoting the commutation, while $[A,B]_{+}=AB+BA$ for the anti-commutation. We denote the characteristic damping rate of the system as $\eta$ and the temperature of heat bath as $T$. Given a quantum operator $O(t)$, its time-dependent expectation value with respect to density matrix $\rho(t)$ is calculated by $\langle O(t)\rangle_{t}=\mathrm{Tr}(\rho(t)O(t))$.
2.1 Quantum Langevin Dynamics (QLD)
In this subsection, we provide a succinct introduction to the genesis of Quantum Langevin Dynamics (QLD) within the Heisenberg picture, and elaborate on our motivation for employing it in optimization tasks. It is well-established that the continuous-time limit of Stochastic Gradient Descent aligns with the classical Langevin equation [71, 69]. Some researches have extended the classical Langevin equation to the quantum case. Ref. [26] put forward the following Hamiltonian for the dynamics of a quantum system coupled to a heat bath:
$H_{\text{total}}=H+\sum_{n}\left(\frac{p_{n}^{2}}{2m_{n}}+\frac{k_{n}}{2}(q_{n}-X)^{2}\right),$ (6)
where $X$ is a specific system operator and the heat bath is simplified as a set of harmonic oscillators with different masses $m_{n}$ and spring constants $k_{n}$ [26, 32]. This kind of coupling characterizes what happens for electromagnetic radiation, but is not appropriate in some other situations such as electrical resistance in a metal [32].
Originated from Eq. (6), Ref. [25] gave a dynamical description of a system moving in a potential, which can be considered as a quantum version of the classical Langevin equation for the motion of system in a potential under the influence of viscous drag and a fluctuating force [32, 25]. Set the Hamiltonian of one-dimensional system to be $H=\frac{p^{2}}{2m}+V(x)$, where $p$ is the system momentum operator, $m$ is the mass of the system, $x$ is the system position operator and $V(x)$ is the potential function. Under the Heisenberg picture, we can get the equation of motion for the system from Eq. (6) by setting $X=x$ [25, 32],
$m\ddot{x}(t)=-V^{\prime}(x(t))-\int_{0}^{t}f(t-t^{\prime})\dot{x}(t^{\prime})\mathrm{d}t^{\prime}-f(t-0)x(0)+\xi(t),$ (7)
where $f(t)$ is the memory function, which makes the equation of motion at time $t$ be influenced by the values of $\dot{x}(t)$ at preceding moments, and $\xi(t)$ is the noise term, with forms
$f(t)=\sum_{n}k_{n}\cos\omega_{n}t,$ (8)
$\xi(t)=\sum_{n}\left(q_{n}(0)k_{n}\cos\omega_{n}t+\dot{q}_{n}(0)\sqrt{k_{n}m_{n}}\sin\omega_{n}t\right).$ (9)
Here, $\omega_{n}=\sqrt{k_{n}/m_{n}}$.
Under the first Markov approximation and the continuum-spectrum approximation [32], we obtain $f(t)=2\gamma\delta(t)$, where $\gamma$ is a damping constant and $\delta(t)$ is the Dirac delta function. Then we get the quantum Langevin equation in the Heisenberg picture, which has the same form as the classical Langevin equation [71, 69, 33] in the underdamping limit but is fully quantum mechanical,
$m\ddot{x}(t)=-V^{\prime}(x(t))-\gamma\dot{x}(t)+\xi(t).$ (10)
While dissipation in quantum mechanics is described by quantum Langevin equation to some extent, analyzing the dynamics of this open quantum system remains to be a such challenging task that we can hardly move a step forward with Eq. (10) in the Heisenberg picture when studying the convergence of QLD. In order to solve this challenge, we introduce the Lindblad functional for QLD in the next section.
2.2 Lindblad functional for QLD
@@ A Lindblad equation can be presented in the standard form:
where $L_{j}$ for $j\in[m]$ are linear opeartors known as jump operators. Ref. [29] studied the case where the Hamiltonian $H$ is given by $H=\frac{p^{2}}{2m}+V$ and the unique Lindblad operator $L$ is a linear combination of the position operator $x$ and the momentum operator $p$,
@@
where $V(x_{1},\ldots,x_{K})\colon\mathbb{R}^{d}\to\mathbb{R}$ is the potential energy (and also the objective function we want to optimize). The Lindblad operator of the $j$-th dimension in space $\mathbb{R}^{d}$ is
@@ 3.1 Convergence of QLD in the convex case
Theorem 1 (Convergence of QLD).
Assume that $V$ is a continuously differentiable convex function and $\nabla^{2}V(x)<+\infty$ for any $x$ in the domain. Let $x^{*}$ be the unique global minimum of $V(x)$. Then the solution $\rho(t)$ to Eq. (14), with approximation $\hbar\Omega\ll kT$, satisfies that
$\left\langle{V}\right\rangle_{t}-V(x^{*})\leq 2kT+O(e^{-\eta t}),$ (21)
if the initial density matrix $\rho(0)$ satisfies $\left\langle{H}\right\rangle_{0}-2kT\geq 0$, where $\langle\cdot\rangle_{t}$ denotes the expectation value with respect to the density matrix $\rho(t)$, i.e., $\langle\cdot\rangle_{t}=\mathrm{Tr}(\cdot\rho(t))$.
Proof.
$\left\langle{V}\right\rangle_{t}\leq 2kT+O(e^{-\eta t}).$ (22)
By Proposition 2, we get $\left\langle{V}\right\rangle_{t}\leq\left\langle{H}\right\rangle_{t}\leq 2kT+O(e^{-\eta t})$. ∎
Remark 1.
If the desired accuracy is $\varepsilon$, which means we want $\left\langle{V}\right\rangle_{t}-V(x^{*})\leq\varepsilon$ when $t\to+\infty$. To achieve this, we only need to set $kT=\varepsilon/2$ and ensure the approximation $\hbar\Omega\ll kT$ still holds.
@@
Let’s focus on the second term $\mathrm{Tr}\left((O(t)\frac{\mathrm{d}\rho(t)}{\mathrm{d}t}\right)$. Inserting Eq. (11) into the second term, we get
@@
Proof.
It is well-known that $[x,p]=\mathrm{i}\hbar$.
$[x,p^{2}]=xp^{2}-p^{2}x=p[x,p]+[x,p]p=2\mathrm{i}\hbar p$
In order to prove Eq. (28b), we introduce a continuously differentiable test function $g(x)$, then
$[f,p]g$ $=\left(f\frac{\hbar}{\mathrm{i}}\nabla-\frac{\hbar}{\mathrm{i}}\nabla f\right)g$
$=\frac{\hbar}{\mathrm{i}}\left(f\nabla g-\nabla fg\right)$
$=\frac{\hbar}{\mathrm{i}}\left(f\nabla g-f\nabla g-g\nabla f\right)$
$=\mathrm{i}\hbar g\nabla f,$
so we get $[f(x),p]=\mathrm{i}\hbar\nabla f$.
Similarly, we have $[\nabla f(x),p]=\mathrm{i}\hbar\nabla^{2}f$.
$[f(x),p^{2}]$ $=f(x)p^{2}-p^{2}f(x)$
$=p[f,p]+[f,p]p$
$=p\cdot\mathrm{i}\hbar\nabla f+\mathrm{i}\hbar\nabla f\cdot p$
$=\mathrm{i}\hbar[\nabla f,p]_{+}.$
$[p,xp]=-[xp,p]=-x[p,p]-[x,p]p=-\mathrm{i}\hbar p.$
Similarly, we can prove $[p,px]=-\mathrm{i}\hbar p$ and $[x,px]=[x,xp]=\mathrm{i}\hbar x$.
$[p^{2},xp]$ $=p[p,xp]+[p,xp]p$
$=-p\cdot\mathrm{i}\hbar p-\mathrm{i}\hbar p\cdot p$
$=-2\mathrm{i}\hbar p^{2}.$
Similarly, we can show that $[p^{2},px]=-2\mathrm{i}\hbar p^{2}$ and $[x^{2},xp]=[x^{2},px]=2\mathrm{i}\hbar x^{2}$. ∎
@@
Assume that $V$ is a continuously differentiable convex function and $\nabla^{2}V(x)<+\infty$ for any $x$ in the domain. If $\hbar\Omega\ll kT$ and the initial condition satisfies $\left\langle{H}\right\rangle_{0}-2kT>0$, the inequality $\left\langle{H}\right\rangle_{t}-2kT\leq(\left\langle{H}\right\rangle_{0}-2kT)e^{-\eta t}=O(e^{-\eta t})$ holds under the evolution of Eq. (14).
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}=$ $\eta\cdot 2kT+\eta\cdot\frac{\hbar\Omega}{kT}\left(\frac{\hbar\Omega}{24}+\frac{\hbar\langle\nabla^{2}V\rangle_{t}}{8m\Omega}\right)+O\left(\left(\frac{\hbar\Omega}{kT}\right)^{3}\right)-\eta\langle x\cdot\nabla V+\frac{p^{2}}{m}\rangle_{t}.$ (34)
The assumption $\nabla^{2}V(x)<+\infty$ and approximation $\frac{\hbar\Omega}{kT}\ll 1$ guarantee that $\eta\cdot\frac{\hbar\Omega}{kT}\cdot\frac{\hbar}{8m\Omega}\langle\nabla^{2}V\rangle_{t}$ is negligible compared to $\eta\cdot 2kT$, which also tells us how small $\frac{\hbar\Omega}{kT}$ must be at least to ensure the approximation work. Neglecting all higher order terms, we get
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}=$ $\eta\cdot\left(2kT-\langle x\cdot\nabla V+p^{2}/m\rangle_{t}\right)$ (35)
$\leq$ $\eta\cdot\left(2kT-\langle x\cdot\nabla V+p^{2}/(2m)\rangle_{t}\right)$
$\leq$ $\eta\cdot\left(2kT-\langle V+p^{2}/(2m)\rangle_{t}\right),$ (36)
where the last inequality comes from the property of convex functions $x\cdot\nabla V(x)\geq V(x)\geq 0$. Finally, we have
$\frac{\mathrm{d}}{\mathrm{d}t}\left\langle{H}\right\rangle_{t}\leq\eta\cdot(2kT-\left\langle{H}\right\rangle_{t}).$ (37)
Construct a function $h(t)=(\left\langle{H}\right\rangle_{t}-2kT)e^{\eta t}$, and it is direct to verify that $\frac{\mathrm{d}h(t)}{\mathrm{d}t}\leq 0$ for $\forall t>0$ by Eq. (37). Consequently, we get $h(t)\leq h(0)$,
$\left\langle{H}\right\rangle_{t}-2kT\leq(\left\langle{H}\right\rangle_{0}-2kT)e^{-\eta t}=O(e^{-\eta t}).$ (38)

### v3

@@ Abstract
We initiate the study of utilizing Quantum Langevin Dynamics (QLD) to solve optimization problems, particularly those nonconvex objective functions that present substantial obstacles for traditional gradient descent algorithms. Specifically, we examine the dynamics of a system coupled with an infinite heat bath. This interaction induces both random quantum noise and a deterministic damping effect to the system, which nudge the system towards a steady state that hovers near the global minimum of objective functions. We theoretically prove the convergence of QLD in convex landscapes, demonstrating that the average energy of the system can converge to zero in the low temperature limit with an exponential convergence rate. Numerically, we first show the energy dissipation capability of QLD by retracing its origins to spontaneous emission. Furthermore, we conduct detailed discussion of the impact of each parameter. Finally, based on the observations when comparing QLD with the classical Fokker-Plank-Smoluchowski equation, we propose a time-dependent QLD by setting temperature and $\hbar$ as time-dependent parameters, which can be theoretically proven to converge better than the time-independent case and also outperforms a series of state-of-the-art quantum and classical optimization algorithms in many nonconvex landscapes.
@@ 1 Introduction
$\dot{Y}=\frac{\mathrm{i}}{\hbar}[H_{\text{total}},Y].$ (2)
Although the above equation intuitively illustrates the Langevin dynamics we consider, it involves the dynamics of each oscillator in the heat bath, making it difficult to capture the energy changes in the system of interest. Therefore, it is necessary to reconsider this problem from a statistical perspective.
We return to the Schrödinger picture and follow the ideas in [47, 8, 31], which adopt the Born-Markov approximation and use path integral approach. When the correlation time of the noise is significantly shorter than the typical time scale of the damped motion of the system, we obtain the quantum Brownian motion master equation:
@@
where $H=p^{2}/(2m)+V$, $p$ is the momentum operator of the system, $x$ is the position operator, and $\eta=\gamma/2m$ represents the characteristic damping rate with system mass $m$, $k$ is the Boltzmann constant. However, it is well known that this master equation is only applicable in high-temperature scenarios and fails to satisfy the positivity constraint of the density operator [28, 19, 43]. Ultimately, we refer to Ref. [28] and employ Lindblad functional (Eq. (9)) to describe the quantum Langevin dynamics, with a detailed introduction provided in Section 2.
@@
where $V(x)$ is assumed to be continuously differentiable in the domain and is also the potential energy of the system in Eq. (9). One can implement Quantum Langevin Dynamics (QLD) by starting from an easily prepared initial state and evolving the quantum system according to Eq. (9). The global minimum of the objective function $V(x)$ can be obtained by measuring the position observable $x$ at the end of the evolution. [^1] Theoretically, the measurement operator is continuous, i.e., |$\int\ket{x}\bra{x}\mathrm{d}x=I$. In practice, the problem is always defined on a finite domain. We discretize the domain of $x$ and the set of eigenstates of $x$ becomes {|$\{\ket{x_{i}}:i=0,1,\ldots,n-1\}$, which can be represented with a set of computational basis. Thus, the measurement operator becomes $\sum_{i}\ket{x_{i}}\bra{x_{i}}$ |. Therefore, our objective is to establish a theoretical foundation guaranteeing the convergence to the global minimum of QLD. When the parameters of QLD is time-independent, namely, $\eta,\hbar$ and $T$ are constants during evolution, we establish the convergence of QLD within convex landscapes (Theorem 1 and 2) and quasar-convex landscape (Theorem 3). Informally speaking, with approximation of $\hbar\Omega\ll kT$ where $\Omega$ is the oscillator frequency of the ground state near the global minimum, we have the convergence rate of the average potential,
$\left\langle{V}\right\rangle_{t}-V(x^{*})\leq\text{$\zeta$}kf(T)+O(e^{-\eta t}),$ (5)
@@
where we denote $\zeta=2+O(\frac{\hbar\Omega}{kT})$ as a constant throughout the paper, $\left\langle{V}\right\rangle_{t}$ denotes the expectation value of $V$ with respect to the density matrix of the system, $V(x^{*})$ is the potential value at the global minimum $x^{*}$ and $f(T)$ is a function with respect to $T$ that satisfies $\lim_{T\to 0}f(T)=0$. Consequently, the average energy of the system can approach zero in the low temperature limit. When $V$ is a quadratic form, the evolution of the system has an analytical solution. One interesting discovery related to the quadratic potential is that the system can reach the ground state of the quadratic potential when the evolution time $t\to+\infty$ and the temperature $T\to 0$. Formal description can be found in Section 3. However, it is always challenging to give a convergence in the general nonconvex case. Despite our best efforts, we are constrained to confine our theoretical discussions to quasar-convex case. Thus, for other nonconvex landscapes, we resort to meticulous numerical experiments to compensate for the absence of theoretical assurances.
@@
In the realm of numerical experiments, we initially substantiate our intuition regarding the fundamental cause of energy dissipation in QLD. Noteworthy is the observation that despite the finite nature of the heat bath, the system retains the capacity to escape local minima, transitioning from excited states to ground states due to spontaneous emission. We conduct experiments within quadratic frameworks, comparing results with analytic solutions to demonstrate strong agreement with our theoretical results. To offer optimal strategies for harnessing thermal and tunneling effects when confronting intricate nonconvex problems, we engage in a detailed discussion delineating the role of each parameter. Broadly speaking, our results indicate that $\eta$ determines the convergence rate, temperature $T$ reflects the thermal effect to some extent and the Plank constant $\hbar$ reflects the tunneling effect.
@@
Building upon these observations, we find it imperative to explore the realm of time-dependent QLD. Specifically, we recognize the need for pronounced thermal and tunneling effects initially to facilitate escape from local minima. As the system converges to its final stage, these effects need to be gradually deactivated. Consequently, we present a theoretical analysis for time-dependent QLD and propose a cooling strategy for implementation. The convergence of time-dependent QLD is similar to the time-independent one when $\hbar=\hbar(t)$ decreasing monotonically with respect to time and $\eta=\eta(t)$, which says $\left\langle{V}\right\rangle_{t}-V(x^{*})\leq\text{$\zeta$}kT+O(e^{-\int_{0}^{t}\eta(t^{\prime})\mathrm{d}t^{\prime}})$. For the case of time-dependent temperature, there exists a cooling strategy $T=T(t)$ that guarantees convergence. In numerical experiments, time-dependent QLD significantly outperforms its time-independent counterpart and surpasses other algorithms, including QHD, QAA, and SGD, across a diverse array of landscapes. In theory, the query complexity of QLD scales as $\tilde{{O}}(t)$, where $\tilde{{O}}(\cdot)$ denotes the asymptotic complexity up to poly-logarithmic factors. This complexity is among the most efficient of all comparable algorithms.”
The rest of this paper is laid out as follows. In Section 2, we introduce our notations and provide details for explicit description of QLD in Eq. (9). Theoretically, we prove the convergence of QLD in the convex case (Theorem 1) and the quasar-convex case (Theorem 3), referring to Section 3. Numerically, we first trace back to spontaneous emission to show the essential cause of energy dissipation in QLD, which verifies our intuition, see Section 4.1. Then we verify our theoretical results of quadratic case in Section 4.3 and discuss the roles of each parameter and their relations with thermal and tunneling effects in Section 4.4. In Section 5 we compare QLD with classical dynamics, which informs the development of time-dependent QLD to fully leverage the advantages inherent in quantum dynamics. In Section 5.1, we prove the convergence of time-dependent QLD in convex case. The performance of time-dependent QLD is evaluated in eight representative landscapes, demonstrating superior performance through comparisons with QHD, QAA, and classical algorithms, shown in Section 5.6. Finally, we evaluate and compare the query complexity of QLD with other algorithms in Section 6, demonstrating that QLD is superior to other algorithms in terms of query complexity.
@@ 2 Preliminaries
Notations. We denote the Hamiltonian of system as $H$, the Hamiltonian of heat bath as $H_{B}$, the Hamiltonian of heat bath and system as $H_{\text{total}}=H+H_{B}$. Let $E_{k}$ be the kinetic energy of the system and $V$ be the potential energy of the system, so $H=E_{k}+V$. Given two quantum operators $A$ and $B$, we use $[A,B]=AB-BA$ for denoting the commutation, while $[A,B]_{+}=AB+BA$ for the anti-commutation. We denote the characteristic damping rate of the system as $\eta$ and the temperature of heat bath as $T$. Given a quantum operator $O(t)$, its time-dependent expectation value with respect to density matrix $\rho(t)$ is calculated by $\langle O(t)\rangle_{t}=\mathrm{Tr}(\rho(t)O(t))$ [^2] We note here that in our setup, the operator $O(t)$ is defined in the Schrödinger picture, which might explicitly depend on time..
@@ A Lindblad equation can be presented in the standard form:
where $L_{j}$ for $j\in[m]$ are linear operators known as jump operators. Ref. [28] studied the case where the Hamiltonian $H$ is given by $H=\frac{p^{2}}{2m}+V$ and the unique Lindblad operator $L$ is a linear combination of the position operator $x$ and the momentum operator $p$,
@@
where $V(x_{1},\ldots,x_{d})\colon\mathbb{R}^{d}\to\mathbb{R}$ is the potential energy (and also the objective function we want to optimize). The Lindblad operator of the $j$-th dimension in space $\mathbb{R}^{d}$ is
@@ 3.1 Convergence of QLD in the convex case
Let us focus on the second term $\mathrm{Tr}\left(O(t)\frac{\mathrm{d}\rho(t)}{\mathrm{d}t}\right)$. Inserting Eq. (6) into the second term, we get
@@
[…]
@@
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}=$ $\eta\cdot 2kT+\eta\cdot\frac{\hbar\Omega}{kT}\left(\frac{\hbar\Omega}{24}+\frac{\hbar\langle\nabla^{2}V\rangle_{t}}{8m\Omega}\right)+O\left(\left(\frac{\hbar\Omega}{kT}\right)^{3}\right)-\eta\left\langle x\cdot\nabla V+\frac{p^{2}}{m}\right\rangle_{t}.$ (26)
[…]
$\frac{\mathrm{d}}{\mathrm{d}t}\langle H\rangle_{t}=$ $\eta\cdot\left(kT\left(2+O\left(\frac{\hbar\Omega}{kT}\right)\right)-\langle x\cdot\nabla V+p^{2}/m\rangle_{t}\right)$ (27)
$\leq$ $\eta\cdot\left(kT\left(2+O\left(\frac{\hbar\Omega}{kT}\right)\right)-\langle x\cdot\nabla V+p^{2}/(2m)\rangle_{t}\right)$
$\leq$ $\eta\cdot\left(kT\left(2+O\left(\frac{\hbar\Omega}{kT}\right)\right)-\langle V+p^{2}/(2m)\rangle_{t}\right),$ (28)
where the last inequality comes from the property of convex functions $x\cdot\nabla V(x)\geq V(x)\geq 0$. Finally, denote $\zeta=2+O\left(\frac{\hbar\Omega}{kT}\right)$, we have
$\frac{\mathrm{d}}{\mathrm{d}t}\left\langle{H}\right\rangle_{t}\leq\eta\cdot(\text{$\zeta$}kT-\left\langle{H}\right\rangle_{t}).$ (29)
Construct a function $h(t)=(\left\langle{H}\right\rangle_{t}-\text{$\zeta$}kT)e^{\eta t}$, and it is direct to verify that $\frac{\mathrm{d}h(t)}{\mathrm{d}t}\leq 0$ for $\forall t>0$ by Eq. (29). Consequently, we get $h(t)\leq h(0)$,
$\left\langle{H}\right\rangle_{t}-\text{$\zeta$}kT\leq(\left\langle{H}\right\rangle_{0}-\text{$\zeta$}kT)e^{-\eta t}=O(e^{-\eta t}).$ (30)
Theorem 1 (Convergence of QLD).
[…]
$\left\langle{V}\right\rangle_{t}-V(x^{*})\leq\text{$\zeta$}kT+O(e^{-\eta t}),$ (31)
if the initial density matrix $\rho(0)$ satisfies $\left\langle{H}\right\rangle_{0}-2kT\geq 0$, where $\zeta$ denotes the constant $2+O\left(\frac{\hbar\Omega}{kT}\right)$, and $\langle\cdot\rangle_{t}$ denotes the expectation value with respect to the density matrix $\rho(t)$, i.e., $\langle\cdot\rangle_{t}=\mathrm{Tr}(\cdot\rho(t))$.
Proof.
$\left\langle{V}\right\rangle_{t}\leq\text{$\zeta$}kT+O(e^{-\eta t}).$ (32)
By Proposition 2, we get $\left\langle{V}\right\rangle_{t}\leq\left\langle{H}\right\rangle_{t}\leq\text{$\zeta$}kT+O(e^{-\eta t})$. ∎
In the theorem above, we prove the convergence of QLD with the assumption $\hbar\Omega\ll kT$. According to the physical significance of these parameters, $\hbar\Omega$ can be seen as a characteristic energy of quantum effect, while $kT$ can be seen as a characteristic energy of thermal effect. Thus, this theorem implies that the convergence in convex cases of QLD is guaranteed by the thermal effect of the heat bath, which provides dissipation. Intuitively, the benefits of quantum effect is reflected in the non-convex situations. Due to the difficulty of theoretical analysis in non-convex case, we verify our intuition with numerical experiments in Section 4.4.
Remark 1.
If the desired accuracy is $\varepsilon$, which means we want $\left\langle{V}\right\rangle_{t}-V(x^{*})\leq\varepsilon$ when $t\to+\infty$. To achieve this, we only need to set $kT=\varepsilon/\text{$\zeta$}$ and ensure the approximation $\hbar\Omega\ll kT$ still holds.